find the value of k if line AB has a slope of -2, when A(-2,-5) and B(k,7)?
Answer by lwsshak3(11628)

find the value of k if line AB has a slope of -2, when A(-2,-5) and B(k,7)?
**
slope=m=-2=∆y/∆x=(y1-y2)/(x1-x2)=(7+5)/(k+2)
-2=12/(k+2)
-2k-4=12
-2k=16
k=-8